# On the Convergence Theory of Pipeline Gradient-based Analog In-memory Training

### v2

@@ III-D Sampling and wall-clock complexity
Sample complexity measures the computational cost required to achieve the desired accuracy, whereas in model parallelism, the computation of different stages overlaps. Therefore, we are also interested in the wall-clock time it takes. In this section, we first study the sample complexity, defined as the minimum number of gradient computations required to achieve the desired error $\epsilon$. Due to asymptotic error, we define the sample complexity in analog training as the minimum micro-batch needed for achieving $E_{K}\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e. $\sigma^{2}S$ for Analog-SGD-WOP/Analog-SGD-SP and $\sigma^{2}S^{\prime}$ for Analog-SGD-AP, respectively. Noticing that the former requires $B$ micro-batches at each iteration while the latter requires $1$, we have the following conclusion.
@@
This shows that three model parallelisms share the same dominant term ${\mathcal{O}}(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while Analog-SGD-AP shows down by a non-dominant term ${\mathcal{O}}(\varepsilon^{-1})$. Since the computing is overlapped, the computation density is also needed to yield the number of clock cycles needed to achieve the $\varepsilon$ accuracy.
@@
CIFAR10
speedup accuracy (mid) accuracy (end)
Digital w.o. pipeline 1.00 88.63  $\pm$ 0.70 94.32  $\pm$ 0.12
synchronous 3.69 94.17  $\pm$ 0.15 94.32  $\pm$ 0.12
asynchronous 6.18 94.25  $\pm$ 0.17 94.25  $\pm$ 0.16
Analog w.o. pipeline 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07
synchronous 3.69 93.75  $\pm$ 0.08 93.95  $\pm$ 0.07
asynchronous 5.66 93.60  $\pm$ 0.05 93.60  $\pm$ 0.07
CIFAR100
speedup accuracy (mid) accuracy (end)
Digital w.o. pipeline 1.00 67.85  $\pm$ 0.52 73.99  $\pm$ 0.19
synchronous 3.69 73.91  $\pm$ 0.19 73.99  $\pm$ 0.19
asynchronous 8.19 74.64  $\pm$ 0.48 74.64  $\pm$ 0.53
Analog w.o. pipeline 1.00 61.35  $\pm$ 0.73 72.29  $\pm$ 0.30
synchronous 3.69 72.14  $\pm$ 0.31 72.29  $\pm$ 0.30
asynchronous 6.17 72.69  $\pm$ 0.30 72.64  $\pm$ 0.21
TABLE II: Training ResNet10 on CIFAR10/CIFAR100 dataset via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). The speedup is computed by comparing the number of clock cycles required to achieve 93% accuracy. The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
@@
TABLE III: Training Resnet10 model on CIFAR10 dataset under different micro-batch size setting. Test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported. The speedup is computed by comparing the number of clock cycles required to achieve 93% target accuracy. The accuracy of synchronous is not affected by micro-batch size.
@@ IV Numerical Simulations
In this section, we aim to numerically verify the theoretical benefits of Analog-SGD-AP. We test the analog methods in an open-source AIHWKIT (AIHWKIT[38]; see github.com/IBM/aihwkit. The pipeline parallelism is also simulated on a single GPU. To get the equivalent clock cycle in the figures, we map the epoch to clock cycle by dividing the computation density.
@@
Simulation settings. In the simulations, we train ResNet models on two benchmark datasets: CIFAR10 and CIFAR100. The CIFAR10 dataset consists of 60,000 $\times32$ color images across 10 classes, with each class representing a distinct category such as airplanes, cars, and birds. The CIFAR100 dataset is similar in format but includes 100 classes. Each model was trained over 300 epochs using a mini-batch size of $B_{\text{mini}}=128$ and micro-batch-size $B_{\text{micro}}=16$ by default. The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage numbers $M=12,16$, the initial learning rate is $\alpha=0.01$. The transfer learning rate of Tiki-Taka is chosen to be equal to the initial learning rate and kept fit during the training. The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs. To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [39]. The other setting follows the default settings of AIHWKIT, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set to 20, output noise 0.1, and input and output quantization to 8 bits). Noise and bound management techniques are used in [40]. A learnable scaling factor is set after each analog layer, which is updated using digital SGD.
@@
Sample complexity and wall-clock complexity. We train a ResNet10 for an image classification task on CIFAR10. We map the first 5 stages on digital accelerators while the last one onto an AIMC (analog) accelerator. The results on the left of Fig. 4 show that the accuracy-epoch curves of all three pipeline strategies on AIMC accelerators almost overlap and even match the performance of digital SGD, implying all of these algorithms have the same sample complexity and align with Theorem 12 and Corollary 1. Moreover, as indicated by Corollary 2, the walk-clock complexity of the Analog-SGD-AP pipeline is the lowest owing to the highest computation density. The results on the CIFAR100 training are reported in Table III, whose conclusion is the same as above.
@@
Effect of stage/accelerator number. We also explored the impact of increasing the number of accelerators on the speedup achieved through pipeline parallelism; see Figure 4, Right. In the simulations, we train a ResNet34 model [41] on the CIFAR10 dataset. The model is separated into $M=1,2,4,8,12,16$ stages. Assuming the model can be evenly divided into multiple stages, and the equivalent clock cycle needed to compute the entire model is fixed. The results reveal that the asynchronous pipeline achieves linear speedup at least within the range of 181-8 accelerators.
@@
Effect of micro-batch size. We also report the effect of micro-batch sizes in Table III. We fit the mini-batch size as $B_{\text{mini}}=128$ and change the micro-batch size. For synchronous pipeline training, increasing the batch size improves the speedup on both digital and analog accelerators. Moreover, asynchronous pipeline training consistently results in higher speedup compared to the synchronous version. The trade-off is that the asynchronous pipeline sacrifices accuracy and can only reach lower accuracy when the training time is sufficiently long.
@@ V Conclusions and Limitations
Appendix
Appendix A Analog SGD with Asynchronous Pipeline in Different Perspective
Involving multiple accelerators and data at each time step, pipeline algorithms, especially the asynchronous pipeline, are usually hard to figure out. To facilitate the understanding, we list the Analog SGD under two different perspectives. In accelerator perspective, we pay more attention to how accelerator $m$ processes a flow of data; see Algorithm 1. In data perspective, we focus on how a sample $\xi_{k}$ flows through each accelerator; see Algorithm 2. To make the relationship clearer, the computation dependency is plotted in Fig. 3.
@@ Appendix B Useful Lemmas and their proof
$\leq$ $\ L_{g,0}((L_{g,0}W_{\max,2})^{m}\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|+W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
where (a) follows the triangle inequality and the consistency of $\ell_{2}$-matrix norm, i.e. $\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrices and vectors, respectively. Using triangle inequality over the first term of RHS in (34), we have
@@
There exists $L_{\ell,0}$ such that for any $m\in[M]$, $\|\nabla_{x^{(m)}}\ell(x_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$ and $\|\nabla_{\tilde{x}^{(m)}}\ell(\tilde{x}_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$.
@@
According to norm equivalence, there exists $W_{\max,2}=cW_{\max}$ with $c$ is defined by the dimension of $W_{k}^{(m)}$, such that $\|W_{k}^{(m)}\|\leq W_{\max,2}$. For $m\in[M]$, according to the forward recursion (3), it holds that
@@
Lemma 8 (boundeds signal).
Under Assumption 24, both forward and backward signals are bounded, i.e., $\forall m\in[M+1]$
@@
This lemma is proved by induction over $m$. The statement for $\|x^{(m)}_{k}\|$ and $\|{\tilde{x}}^{(m)}_{k}\|$ at $m=1$ trivially hold because $x^{(0)}_{k}={\tilde{x}}^{(0)}_{k}=x_{k}$. For $\|\delta^{(M+1)}_{k}\|$ and $\|\tilde{\delta}^{(M+1)}_{k}\|$, we have
@@
$\leq$ $\ \|\delta^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}\|$,
@@
$\leq$ $\ \|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}\|\tilde{\delta}^{(m+1)}_{k}\|$.
The proof has finished. ∎

### v3

@@ III-D Sampling and wall-clock complexity
Sample complexity measures the computational cost required to achieve the desired accuracy, whereas in model parallelism, the computation of different stages overlaps. Therefore, we are also interested in the wall-clock time it takes. In this section, we first study the sample complexity, defined as the minimum number of gradient computations required to achieve the desired error $\varepsilon$. Due to asymptotic error, we define the sample complexity in analog training as the minimum micro-batch needed for achieving $E_{K}\leq\varepsilon+E$, where $E$ is the asymptotic error, i.e. $\sigma^{2}S$ for Analog-SGD-WOP/Analog-SGD-SP and $\sigma^{2}S^{\prime}$ for Analog-SGD-AP, respectively. Noticing that the former requires $B$ micro-batches at each iteration while the latter requires $1$, we have the following conclusion.
@@
This shows that three model parallelisms share the same dominant term ${\mathcal{O}}(\sigma^{2}\varepsilon^{-2})$ in the sample complexity, while Analog-SGD-AP shows down by a non-dominant term ${\mathcal{O}}(\varepsilon^{-1})$. Since the computation is overlapped, the computation density is also needed to determine the number of clock cycles required to achieve $\varepsilon$ accuracy.
@@
Remark 6 (Load balancing in pipeline parallelism).
The computation density analysis above assumes that the DNN model is evenly partitioned across $M$ pipeline stages, such that each stage incurs the same computational cost per micro-batch. In practice, uneven partitioning would introduce pipeline bubbles at the bottleneck stage, reducing the effective computation density. However, this load-balancing issue is orthogonal to the convergence analysis of this paper. Theorems 1 and 2, as well as Corollary 1, remain valid regardless of how the model is partitioned, as they bound the sample complexity rather than the wall-clock complexity directly. For Corollary 2, a computation density of 1 is an achievable upper bound under balanced partitioning. If the partition is unbalanced, the density should be scaled down by a partition-specific factor, which approaches 1 as partitioning becomes more balanced. In practice, balanced partitioning can be achieved using existing pipeline-parallelism frameworks such as [huang2019gpipe, narayanan2019pipedream, shoeybi2019megatron]; a detailed investigation of load-balancing strategies for AIMC pipelines is beyond the scope of this paper.
Remark 7 (Comparison with digital-SGD pipelines).
In fact, the digital-SGD counterpart is a special case of (1) by forcing $\tau\to\infty$, which removes the asymmetric bias term. In particular, the analog-specific denominator inflation, $\frac{1}{1-W_{\max,\infty}^{2}/\tau^{2}}$ and $\frac{1}{1-(1+u)W_{\max,\infty}^{2}/\tau^{2}}$, and asymptotic error terms $S$ and $S^{\prime}$ vanish in the digital limit. Consequently, the pipelining benefit established in Corollaries 1 and 2 is not unique to AIMC dynamics and can be reproduced in the digital domain.
CIFAR10 CIFAR100
speedup accuracy (mid) accuracy (end) speedup accuracy (mid) accuracy (end)
Digital w.o. pipeline 1.00 88.63  $\pm$ 0.70 94.32  $\pm$ 0.12 1.00 67.85  $\pm$ 0.52 73.99  $\pm$ 0.19
synchronous 3.69 94.17  $\pm$ 0.15 94.32  $\pm$ 0.12 3.69 73.91  $\pm$ 0.19 73.99  $\pm$ 0.19
asynchronous 6.18 94.25  $\pm$ 0.17 94.25  $\pm$ 0.16 8.19 74.64  $\pm$ 0.48 74.64  $\pm$ 0.53
Analog w.o. pipeline 1.00 88.04  $\pm$ 0.58 93.95  $\pm$ 0.07 1.00 61.35  $\pm$ 0.73 72.29  $\pm$ 0.30
synchronous 3.69 93.75  $\pm$ 0.08 93.95  $\pm$ 0.07 3.69 72.14  $\pm$ 0.31 72.29  $\pm$ 0.30
asynchronous 5.66 93.60  $\pm$ 0.05 93.60  $\pm$ 0.07 6.17 72.69  $\pm$ 0.30 72.64  $\pm$ 0.21
TABLE II: ResNet10 @ CIFAR10/CIFAR100 via vanilla model parallelism without pipeline (wo pipeline), synchronous pipeline (Sync), and asynchronous pipeline (Asyn). The speedup is computed by comparing the number of clock cycles required to achieve 93% accuracy. The test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported.
@@
TABLE III: ResNet10 @ CIFAR10 under different micro-batch size setting. Test accuracy and standard deviation at clock cycle 300 (mid) and epoch 300 (end) are reported. The speedup is computed by comparing the number of clock cycles required to achieve 93% target accuracy. The accuracy of synchronous is not affected by micro-batch size.
@@ IV Numerical Simulations
In this section, we aim to numerically verify the theoretical benefits of Analog-SGD-AP. We test the analog methods in an open-source AIHWKIT (AIHWKIT[Rasch2021AFA]; see github.com/IBM/aihwkit. The pipeline parallelism is also simulated on a single GPU. To get the equivalent clock cycle in the figures, we map the epoch to clock cycle by dividing the computation density.
@@
Simulation settings. In the simulations, we train ResNet models on two benchmark datasets: CIFAR10 and CIFAR100. The CIFAR10 dataset consists of 60,000 $\times32$ color images across 10 classes, with each class representing a distinct category such as airplanes, cars, and birds. The CIFAR100 dataset is similar in format but includes 100 classes. Each model was trained over 300 epochs using a mini-batch size of $B_{\text{mini}}=128$ and micro-batch-size $B_{\text{micro}}=16$ by default. The initial learning rate of $\alpha=0.1$ was employed, while for Resnet34 with stage numbers $M=12,16$, the initial learning rate is $\alpha=0.01$. The learning rate $\alpha$ was reduced by a factor of 10 every 100 epochs. To further enhance the model’s ability to generalize across diverse images, we applied data augmentation techniques such as random cropping, horizontal flipping, and color jittering [cubuk2018autoaugment].
AIMC hardware configurations. The simulations use the SoftBoundsReference device in AIHWKIT, which directly instantiates the asymmetric update model in (6) with $\tau=0.9$ by default. Beyond the asymmetric update, the simulations additionally incorporate a comprehensive set of hardware imperfections to more accurately reflect real device behavior, including output noise (0.5 % of the quantization bin width), quantization and clipping (output range set to 20, output noise 0.1, and input and output quantization to 8 bits). In addition, 30% device and cycle variation are introduced. Noise and bound management techniques are used in [gokmen2017cnn]. A learnable scaling factor is applied after each analog layer and updated via digital SGD.
@@
Sample complexity and wall-clock complexity. We train a ResNet10 on CIFAR10. We map the first 5 stages on digital accelerators while the last one onto an AIMC (analog) accelerator. The results on the left of Fig. 3 show that the final accuracy differences between the three strategies are below 1%, implying all of these algorithms have the same sample complexity and align with Theorem 12 and Corollary 1. Moreover, as indicated by Corollary 2, the walk-clock complexity of the Analog-SGD-AP pipeline is the lowest owing to the highest computation density. The results on the CIFAR100 training are reported in Table III, whose conclusion is the same as above.
@@
Effect of stage/accelerator number. We also explored the impact of increasing the number of accelerators on the speedup achieved through pipeline parallelism; see Fig. 3, Right. In the simulations, we train a ResNet34 model [wu2019wider] on the CIFAR10 dataset. The model is separated into $M=1,2,4,8,12,16$ stages. Assuming the model can be evenly divided into multiple stages, and the equivalent clock cycle needed to compute the entire model is fixed. The results reveal that the asynchronous pipeline achieves linear speedup at least within the range of 181-8 accelerators.
As the $M$ increases, the speedup deviates from linear scaling, which can be explained by Theorem 2. As shown by (2), the iteration complexity of Analog-SGD-AP contains a higher-order term $O(M^{3}(1+1/u)/K)$, whose coefficient depends on $M$. Although this term is higher-order and does not affect the dominant sample complexity $O(\sigma^{2}\varepsilon^{-2})$, it has a non-negligible practical impact when $M$ increases while iterations $K$ remain small. Consequently, it results in the observed deviation from linear speedup. Developing acceleration strategies to mitigate this depth-dependent overhead and recover linear speedup at larger scales ($M>8$) is an interesting direction we leave for future work.
@@
Effect of micro-batch size. We also report the effect of micro-batch sizes in Table III. We set the mini-batch size to $B_{\text{mini}}=128$ and vary the micro-batch size. For synchronous pipeline, increasing the batch size improves the speedup on both digital and analog accelerators. Moreover, asynchronous pipeline training consistently yields greater speedup than the synchronous version. The trade-off is that the asynchronous pipeline sacrifices accuracy and can only reach lower accuracy when the training time is sufficiently long.
Effect of hardware imperfections. To disentangle the contributions of analog imperfections and asynchronous pipeline staleness to the convergence error, we compare the performance of digital and analog pipelines under identical configurations, as shown in Tables III and III. The results in Table III show that (i) the accuracy gap between digital and analog is consistent ($1\%-2\%$) across all three pipeline strategies; (ii) within each hardware setting (digital or analog), the accuracy difference between synchronous and asynchronous pipelines is negligible at convergence ($1\%$). These results indicate that pipeline staleness introduces only marginal additional error, and the gap is attributable to analog imperfections. Table III further shows that this conclusion holds robustly across different micro-batch sizes: the accuracy gap between digital and analog remains stable regardless of micro-batch size, while the gap between synchronous and asynchronous pipelines remains small in both settings. Taken together, these results confirm that analog imperfection is the dominant source of convergence error, while the asynchronous pipeline staleness has a secondary effect.
Effect of asymmetry parameter $\tau$. We further examine how the asymmetry parameter $\tau$ affects both convergence and acceleration; see Fig. 4a/b. The left figure shows that when $\tau>0.5$, Analog-SGD-AP still achieves above 90% test accuracy. Furthermore, as shown in the right figure, Analog-SGD-AP maintains a speedup well above the synchronous pipeline baseline (shown as the dashed line at $3.69\times$) for all $\tau\geq 0.4$. This confirms that the wall-clock gains of the asynchronous pipeline are preserved across a wide range of $\tau$ values. In contrast, as $\tau$ decreases, training quality degrades significantly, with asynchronous training affected more strongly.
Figure 4: Analog-SGD-AP for CIFAR10 with $B_{\text{micro}}=16$ and delay =0=0. (a) Test accuracy versus $\tau$. (b) Speedup versus $\tau$. (c) Speedup versus device-to-device variation. (d) Saturation degree evolution over the trajectory.
Effect of device variation on speedup. We next evaluate whether device-to-device variation changes the acceleration trend; see Fig. 4c. In the simulation, we set $\tau=1.0$, $B_{\text{micro}}=16$, and delay =0=0. Across the typical variation levels tested here, the asynchronous pipeline speedup changes only mildly, suggesting that device variation is not the primary limiter of pipeline acceleration in this regime.
Saturation degree evolution along training trajectory. We examine the saturation degree over the full training trajectory to justify the bounded saturation condition; see Fig. 4d. The saturation degree of $m$-th layer is defined as $\|W_{k}^{(m)}\|_{\infty}/\tau$. To ensure the formulation is well-defined, we set the device variation to 0. The results show that the saturation remains nearly constant throughout training, with variations less than 10%. This observation is consistent with existing results, such as [zou2019improved], which demonstrate that the trajectory is bounded in a neighborhood of the initial point. It provides empirical evidence for the Assumption 5 in the studied configuration; a general theoretical characterization is left for future work.
@@ V Conclusions and Limitations
Supplementary
Appendix A Analog SGD with Asynchronous Pipeline in Different Perspectives
Involving multiple accelerators and data at each time step, pipeline algorithms, especially the asynchronous pipeline, are usually hard to figure out. To facilitate the understanding, we list the Analog SGD under two different perspectives. In accelerator perspective, we pay more attention to how accelerator $m$ processes a flow of data; see Algorithm 1. In data perspective, we focus on how a sample $\xi_{k}$ flows through each accelerator; see Algorithm 2. To clarify the relationship, the computational dependency is plotted in Fig. 5.
@@ Appendix B Useful Lemmas and their proof
$\overset{(b)}{\leq}$ $\ L_{g,0}((L_{g,0}W_{\max,2})^{m}\|W^{(m)}_{k-(M-m)}-W^{(m)}_{k}\|+W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|)$
where (a) follows the triangle inequality and the consistency of $\ell_{2}$-matrix norm, i.e. $\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrices and vectors, respectively; and (b) is because of Lemma 8. Using triangle inequality over the first term of RHS in (35), we have
@@
Under same conditions with Lemma 8, there exists $L_{\ell,0}$ such that for any $m\in[M]$, $\|\nabla_{x^{(m)}}\ell(x_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$ and $\|\nabla_{\tilde{x}^{(m)}}\ell(\tilde{x}_{k}^{(m)},y_{k})\|\leq L_{\ell,0}$.
@@
According to norm equivalence, there exists $W_{\max,2}=cW_{\max}$ with $c$ is defined by the dimension of $W_{k}^{(m)}$, such that $\|W_{k}^{(m)}\|\leq W_{\max,2}\leq\tau$. For $m\in[M]$, according to the forward recursion (3), it holds that
@@
Lemma 8 (Bounded signal and iterates).
Under Assumption 24, selecting stepsize $\alpha\leq\frac{\tau}{(L_{g,0}\tau)^{M}L_{\ell,0}L_{g,0}}$ and initializing $\|W_{0}^{(m)}\|<\tau,\forall m\in[M+1]$, then both forward and backward signals are bounded, i.e., $\forall m\in[M+1]$
@@
Moreover, for any $k\leq K$ and $m\in[M+1]$, the iterates $\|W_{k}^{(m)}\|<\tau$ are bounded within the dynamic range, i.e. $W_{\max,2}:=\max_{k\leq K,m\in[M+1]}\{\|W_{k}^{(m)}\|\}<\tau$.
@@
This lemma is proved by induction over $m$ and $k$. We can initialize such that $\|W_{0}^{(m)}\|<\tau$ holds for any $m\in[M+1]$.
At each iteration $k$, assume $\|W_{k}^{(m)}\|<\tau$ holds for any $m\in[M+1]$ and denote $W_{\max,2}^{(m)}=\max_{m}\{\|W_{k}^{(m)}\|\}$. Clearly, $W_{\max,2}^{(m)}<\tau$. Then the statement for $\|x^{(m)}_{k}\|$ and $\|{\tilde{x}}^{(m)}_{k}\|$ at $m=1$ trivially hold because $x^{(0)}_{k}={\tilde{x}}^{(0)}_{k}=x_{k}$. For $\|\delta^{(M+1)}_{k}\|$ and $\|\tilde{\delta}^{(M+1)}_{k}\|$, we have
@@
$\leq$ $\ \|\delta^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}(z^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}^{(m)}\|\delta^{(m+1)}_{k}\|$,
@@
$\leq$ $\ \|\tilde{\delta}^{(m+1)}_{k}\|\|W^{(m+1)}_{k}\|\|\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|\leq L_{g,0}W_{\max,2}^{(m)}\|\tilde{\delta}^{(m+1)}_{k}\|$.
Therefore, by induction over $m$, we obtain,
$\|x^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{m},$ $\|\delta^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{M-m}L_{\ell,0}L_{g,0},$ (61)
$\|{\tilde{x}}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{m},$ $\|\tilde{\delta}^{(m)}_{k}\|\leq$ $\ (L_{g,0}W_{\max,2}^{(m)})^{M-m}L_{\ell,0}L_{g,0}.$ (62)
On the other hand, if we select the stepsize $\alpha$ such that $\alpha\leq\frac{\tau}{(L_{g,0}\tau)^{M}L_{\ell,0}L_{g,0}}$, we have the magnitude of each update increment
$\|\Delta\tilde{W}_{k}^{(m)}\|=\alpha\|{\tilde{x}}^{(m)}_{k}\|\|\tilde{\delta}^{(m)}_{k}\|\leq\tau.$
Owing to the induction nature proof for Theorem 1 (see [wu2024towards, Theorem 1]), we know $\|W_{k+1}^{(m)}\|<\tau$ holds for any $m\in[M+1]$.
By induction, the proof is completed. ∎